A firm demand curve is P =1-2Q. The firm has a current price of R1000 and it sells 100 units per day. What is the firms price elasticity of demand?


1
Expert's answer
2022-04-17T17:03:21-0400

P = 1 - 2Q, so Q = 0.5 - 0.5P.

The firms price elasticity of demand is:

"Ed = -0.5\u00d71000\/100 = -5."

So, the demand is elastic at this price.